# FoMEMO: Towards Foundation Models for Expensive Multi-objective Optimization

### v1

@@ 4.2 Data Generation
Given that expensive data of real-world experiments are typically scarce and users’ potential preferences can be difficult to obtain, we implement a data generation mechanism designed to simulate a substantial amount of potential scenarios that the model may encounter. This process enables us to create a diverse and lightweight dataset that enhances the model’s ability to generalize and respond rapidly to various situations, ultimately improving its optimization performance in emerging real-world tasks.
Our data generation mechanism consists of three steps. Firstly, we sample general parameters for each set of the data, including the number of features $d$, the number of objectives $m$ and the length of trajectory $n$. Secondly, in order to simulate the multitude of implicit mappings that may exist in the real world, we sample a variety of synthetic functions from Gaussian processes (GPs). The choice of GPs is motivated by their cost-effectiveness and flexibility. Sampling from a GP prior is a easy and cost-effective method that enables the generation of a substantial number of latent mappings at a low cost. Besides, we can easily control the diversity of the generated functions by adjusting the kernel parameters. In the third step, we sample the observed trajectory $D_{n}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n}$ together with an arbitrary query pair $(\boldsymbol{x},\boldsymbol{y})$ with $d$-dimensional features and $m$-dimensional objectives from each sampled function. Since the aggregation target $g$ is determined only by $\boldsymbol{x}$ given $\boldsymbol{\lambda}$, after sampling artificial preferences from the simplex $\Delta=\{\boldsymbol{\lambda}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{\lambda}\|_{1}=1\}$, we can calculate the aggregation targets efficiently for each query input given arbitrary preferences. We synthesize hundreds of millions of training data using the aforementioned method, and mask the query aggregation targets to formulate a supervised prediction problem. This processing allows the foundation model to learn to effectively infer the preference-wise posterior distribution $q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})$.
@@ 4.3 Model Architecture
Among the model architectures capable of learning uncertainty-calibrated predictions directly from data, we utilize the Prior-Data Fitted Networks (PFNs) (Müller et al. 2021) as the base framework, as it provides a flexible paradigm for approximating the posterior distribution with arbitrary shape in a principled manner (Nagler 2023). PFNs are based on a Transformer encoder that does not incorporate positional encodings, allowing the model to process data regardless of the order in which the input elements are presented. We use a variable-dimensional encoder to encode each input $\boldsymbol{x}$ and observation $\boldsymbol{y}$ into a token, allowing the model to adapt to cross-dimensional feature and objective spaces. An attention mask is implemented to enable each trajectory point can only attend to each other and the query point can only attend to the trajectory points. We also encode the sampled preference to inform the model to learn to predict the posterior distribution conditioned on which preference. To approximate the complex posterior distribution in high-dimensional feature and objective spaces, PFNs employ a piecewise constant distribution in the regression head. This approach converts the continuous distribution into a discrete distribution by utilizing multiple intervals, thereby transforming the regression task into a classification task to enhance the robustness during training. The detailed implementation can be referred to (Müller et al. 2021).
@@ 4.4 Training Details
$\mathbb{E}_{(\boldsymbol{x},y)\cup D_{n}\sim p(D),\boldsymbol{\lambda}\sim\Delta,g=-s_{\boldsymbol{\lambda}}(\boldsymbol{x})}[-\log q_{\theta}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})].$ (5)
@@
Our final foundation model uses the PFNs architecture with 12 layers of Transformer, an embedding size of 512, a hidden size 1024 in feed-forward layers, and 4-head attention, resulting a total of 26.81M parameters. During training, we randomly sample each dataset from a large data space with feature dimensions ranging from 1 to 30, objective dimensions from 1 to 6, and trajectory lengths from 1 to 128. We sample a large number of functions from a standard RBF-kernel GP with zero mean. We randomly sample the length scale kernel parameter $l\sim\Gamma(\alpha=3.0,\beta=6.0)$ for each feature and add a random observation noise $\epsilon\sim\mathcal{N}(0,10^{-4})$ to ensure function diversity. We do not sample the output scale parameter but instead set it to 1.0, as the predicted aggregation target will be normalized during inference. The final model is trained using 500 epochs and 2048 steps in each epoch with a batch size of 256 datasets, resulting more than 200 million datasets in total. The loss function is optimized using Adam optimizer (Kingma and Ba 2014) with linear warmup and cosine annealing (Loshchilov and Hutter 2016) using a learning rate of $10^{-4}$. The entire training process is conducted on a single NVIDIA A100 GPU and takes approximately 140 hours. We performed the pre-training only once and used the same model in all experiments in this paper.
@@ 4.5 Generating Candidates
Benefiting from the preference-wise aggregation posterior as the output basis, modeling the distribution of infinite sub-problems becomes possible. Acquisition functions such as random scalarization and hypervolume aggregation can be well adapted directly. We have developed two types of acquisition functions including preference-based and preference-free in our work.
We firstly derive the preference-based expected improvement (EI) and upper confidence bound (UCB) acquisition functions based on the single-objective scenario (Müller et al. 2023):
$\alpha_{\text{EI}}(\boldsymbol{x};\boldsymbol{\lambda})=\mathbb{E}_{q_{\theta}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})}[[g_{\boldsymbol{\lambda}}(\boldsymbol{x})-g_{\boldsymbol{\lambda}}^{*}]_{+}],$ (6)
@@
where $g_{\boldsymbol{\lambda}}^{*}$ is the best value of the aggregation $g_{\boldsymbol{\lambda}}(\boldsymbol{x})$ given an arbitrary preference $\boldsymbol{\lambda}$. ${\mu}_{\boldsymbol{\lambda}}(\boldsymbol{x})$ and ${\sigma}_{\boldsymbol{\lambda}}(\boldsymbol{x})$ are the mean and standard deviation of the aggregated posterior $q_{\theta}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})$, respectively. $\beta$ is a constant, which we set to 1.0 in this paper. As inference the acquisition functions is conditioned on the preference, during the optimization process, we randomly sample the preferences in each iteration and optimize the acquisition function to generate the next candidate solutions for evaluation.
To marginalize the preferences, we draw inspiration from a key observation: using infinite reference vectors uniformly sampled from the unit sphere $S=\{\boldsymbol{w}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{w}\|_{2}=1\}$, the hypervolume metric can be represented as the mean of a scalarization function (Deng and Zhang 2019; Zhang and Golovin 2020). Based on this insight, we develop a preference-free acquisition function that uses the UCB utility in Equality (7) to approximate the hypervolume improvment with uncertainty (UHVI) for any $\boldsymbol{x}$ given the evaluated inputs $\boldsymbol{X}=\{\boldsymbol{x}_{i}\}_{i=1}^{n}$:
@@
where $c_{m}=\frac{\pi^{m/2}}{2^{m}\Gamma(m/2+1)}$ is a constant that depends only on $m$. $c_{\boldsymbol{\lambda}}=\sqrt{\sum_{j=1}^{m}\frac{1}{\lambda_{j}^{2}}}$ is a transformation constant that depends only on $\boldsymbol{\lambda}$. It enables us to convert sampling reference vectors from the unit sphere to sampling preferences from the simplex, and can be adapted to our model by simply using the transformation $w_{j}=1/({\lambda_{j}c_{\boldsymbol{\lambda}})}$. The detailed derivation can be found in the appendix.
@@ 5 Experiments
We compare FoMEMO against existing state-of-the-art model-based methods including qParEGO (Knowles 2006; Daulton, Balandat, and Bakshy 2020), qEVHI (Daulton, Balandat, and Bakshy 2020) and qNEHVI (Daulton, Balandat, and Bakshy 2021), all of which demonstrate promising scalability in scenarios involving noise interruptions and parallel optimization. Additionally, we include a Sobol sequence-based random sampling method as a trivial baseline for comparison.
@@
We empirically evaluate the algorithms on various synthetic benchmarks and real-world applications, encompassing a diverse set of feature and objective spaces, as well as various Pareto fronts characteristics with convex, concave and disconnected shapes. For synthetic benchmarks, we consider the popular ZDT (Zitzler, Deb, and Thiele 2000) and Omnitest (Deb and Tiwari 2008) test suits. For real-world applications, we use the RE test suite (Tanabe and Ishibuchi 2020) consisting various multi-objective engineering design problems, such as four bar truss design, reinforced concrete beam design, rocket injector design, etc. The detailed descriptions can be referred to appendix.
Figure 2: Sequential optimization performance of IGD+ metrics ($\downarrow$) w.r.t. the number of observations (beyond initial points) on synthetic problems. The solid line is the mean value averaged over 10 independent runs, and the shaded region is the standard deviation with 95% confidence around the mean value.
@@
For model-based algorithms, we run each experiment using $2(d+1)$ initial samples and a total of 100 evaluations beyond the initial points. The implementations of qEVHI, qNEHVI and qParEGO are based on BoTorch (Balandat et al. 2020), we use the default setting that optimizes the acquisition functions via multi-start L-BFGS-B optimizer with 20 restarts seeded from 1024 pseudo-random samples through BoTorch’s initialization heuristic, we keep the same parameter setting in FoMEMO.
Table 1: Means (stds) of best hypervolume metrics ($\uparrow$) in sequential optimization on real-world problems over 10 independent runs. The best and second best mean results for each row are highlighted with shaded background and underlines respectively.
Test Instances Sobol qParEGO qEHVI qNEHVI FoMEMO-EI FoMEMO-UCB FoMEMO-UHVI
RE21 7.366e-01(1.22e-02) 8.586e-01(6.61e-03) 8.825e-01(1.11e-04) 8.824e-01(2.66e-04) 8.683e-01(1.62e-03) 8.720e-01(1.37e-03) 8.629e-01(1.06e-03)
RE22 5.376e-01(2.88e-02) 5.473e-01(3.81e-02) 5.333e-01(4.46e-02) 5.732e-01(3.80e-02) 4.982e-01(4.31e-02) 5.043e-01(7.61e-02) 6.242e-01(3.42e-02)
RE23 4.107e-01(1.18e-01) 1.111e+00(9.70e-03) 1.119e+00(8.77e-03) 1.120e+00(9.26e-03) 1.154e+00(3.32e-03) 1.156e+00(1.04e-03) 1.011e+00(5.45e-02)
RE24 9.695e-01(1.19e-01) 1.164e+00(1.29e-03) 1.138e+00(4.64e-03) 1.153e+00(3.83e-03) 1.164e+00(1.74e-03) 1.161e+00(1.20e-03) 1.123e+00(8.61e-03)
RE25 1.059e-01(9.92e-02) 9.064e-01(5.54e-02) 9.932e-01(2.13e-02) 9.531e-01(2.01e-02) 1.053e+00(1.32e-02) 1.047e+00(1.02e-02) 6.843e-01(7.65e-02)
RE31 1.233e+00(3.56e-02) 1.318e+00(4.60e-03) 1.270e+00(1.33e-02) 1.300e+00(5.35e-03) 1.314e+00(3.42e-03) 1.310e+00(4.70e-03) 1.298e+00(1.97e-02)
RE32 1.273e+00(1.30e-02) 1.316e+00(6.28e-03) 1.321e+00(3.85e-03) 1.318e+00(4.94e-03) 1.326e+00(1.31e-03) 1.327e+00(1.83e-03) 1.294e+00(8.23e-03)
RE33 1.130e+00(3.50e-02) 1.293e+00(7.83e-03) 1.305e+00(5.02e-03) 1.310e+00(1.01e-03) 1.310e+00(6.72e-04) 1.312e+00(5.31e-04) 1.272e+00(6.93e-03)
RE34 5.788e-01(2.85e-02) 9.183e-01(1.40e-02) 1.039e+00(4.80e-04) 1.038e+00(8.30e-04) 9.890e-01(6.40e-03) 1.005e+00(8.92e-03) 9.367e-01(1.00e-02)
RE35 1.186e+00(9.18e-03) 1.266e+00(8.01e-03) 1.294e+00(1.73e-03) 1.294e+00(2.16e-03) 1.294e+00(3.86e-03) 1.297e+00(9.41e-04) 1.260e+00(6.02e-03)
RE36 3.135e-01(4.38e-02) 8.161e-01(2.77e-02) 8.877e-01(8.70e-03) 8.958e-01(7.11e-03) 9.175e-01(6.87e-03) 9.192e-01(4.87e-03) 8.566e-01(1.13e-02)
RE37 6.046e-01(2.31e-02) 8.068e-01(1.97e-02) 8.862e-01(8.58e-04) 8.864e-01(1.07e-03) 8.515e-01(3.69e-03) 8.642e-01(2.99e-03) 8.325e-01(4.08e-03)
Figure 3: Parallel optimization performance (q=5) of hypervolume metrics ($\uparrow$) w.r.t. the number of observations (beyond initial points) on real-world applications. The solid line is the mean value averaged over 10 independent runs, and the shaded region is the standard deviation with 95% confidence around the mean value.
Performance Metric
To compare the performance of different algorithms on synthetic problems with known true Pareto fronts, we utilize the inverted generational distance plus (IGD+) indicator (Ishibuchi et al. 2015) as the metric to evaluate the quality of the final solutions obtained by each algorithm. For real-world applications, we use the hypervolume indicator (Zitzler and Thiele 2002) as the metric in the normalized objective space for fair comparison. Each experiment is conducted with 10 independent runs. See appendix for more detailed experimental settings.
@@ 5.1 Optimization Performance
Figure 2 shows the sequential optimization performance of IGD+ metrics w.r.t. the number of observations on synthetic problems. The experimental results indicate that the three FoMEMO methods are highly competitive when compared to existing approaches. Specifically, FoMEMO-UCB has demonstrated superior performance compared to other algorithms in most of the tested problems.
Table 1 represents the results of sequential optimization on real-world applications in terms of the hypervolume metric. Experimental results indicate that our method together with the three proposed acquisition functions achieve superior performance in most cases. This demonstrates that by pre-training the foundation model with a large set of diverse synthetic data, the model becomes highly adaptable and generalizable to complex problems that it has not encountered during training, without having to access to large amounts of expensive real-world experimental data.

### v2

@@ 4.2 Data Generation
While the pre-training distribution should ideally align with that of real-world functions, capturing such a distribution is often challenging due to limited empirical data. To bridge this gap, we implement a data generation mechanism designed to simulate a vast array of surrogate scenarios that the model may encounter. This approach enables the creation of a diverse yet computationally efficient dataset, bolstering the model’s generalization capacity and enabling rapid adaptation to complex, unseen real-world optimization tasks.
Our data generation mechanism consists of three steps. Firstly, we sample general parameters for each set of the data, including the number of features $d$, the number of objectives $m$ and the length of trajectory $n$. Secondly, in order to simulate the multitude of implicit mappings that may exist in the real world, we sample a variety of synthetic functions from Gaussian processes (GPs). The choice of GPs is motivated by their cost-effectiveness and flexibility. Sampling from a GP prior is an easy and cost-effective method that enables the generation of a substantial number of latent mappings at a low cost. Besides, we can easily control the diversity of the generated functions by adjusting the kernel parameters. We employ a GP with the RBF kernel as it is a universal approximator (Micchelli et al., 2006) to encompass a wide range of unknown functions, which means in theory if trained long enough the foundation model will encounter synthetic functions that closely resemble true objective functions, enabling effective generalization. In the third step, we sample the observed trajectory $D_{n}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n}$ together with an arbitrary query pair $(\boldsymbol{x},\boldsymbol{y})$ with $d$-dimensional features and $m$-dimensional objectives from each sampled function. Since the aggregation target $g$ is determined only by $\boldsymbol{x}$ given $\boldsymbol{\lambda}$, after sampling artificial preferences from the simplex $\Delta=\{\boldsymbol{\lambda}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{\lambda}\|_{1}=1\}$, we can calculate the aggregation targets efficiently for each query input given arbitrary preferences. We synthesize hundreds of millions of training data using the aforementioned method, and mask the query aggregation targets to formulate a supervised prediction problem. This processing allows the foundation model to learn to effectively predict the preference-wise posterior distribution $q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})$.
@@ 4.3 Model Architecture
Among the model architectures capable of learning uncertainty-calibrated predictions directly from data, we utilize the Prior-Data Fitted Networks (PFNs) (Müller et al., 2021) as the base framework, as it provides a flexible paradigm for approximating the posterior distribution with arbitrary shape in a principled manner (Nagler, 2023). PFNs are based on a Transformer encoder that does not incorporate positional encodings, allowing the model to process data regardless of the order in which the input elements are presented. We use a variable-dimensional encoder to encode each input $\boldsymbol{x}$ and observation $\boldsymbol{y}$ into a token, allowing the model to adapt to cross-dimensional feature and objective spaces. An attention mask is implemented to enable each trajectory point can only attend to each other and the query point can only attend to the trajectory points. We also encode the sampled preference to inform the model to learn to predict the posterior distribution conditioned on which preference. To approximate the complex posterior distribution in high-dimensional feature and objective spaces, PFNs employ a piecewise constant distribution in the regression head. This approach converts the continuous distribution into a discrete distribution by utilizing multiple intervals, thereby transforming the regression task into a classification task to enhance the robustness during training. The detailed implementation can be referred to (Müller et al., 2021). To account for the distributional shifts in aggregated posteriors across varying objective dimensions, we propose an objective-aware regression head. This architecture dynamically routes the input to a specialized regressor tailored to the specific objective dimension of the problem at hand. By leveraging this dimension-specific mapping, the model can more accurately capture the differences of the aggregated posteriors with higher fidelity, thereby significantly enhancing its predictive precision in unseen landscapes across varying objective dimensions. Please refer to Appendices A.1 and C for detailed model configurations and the ablation analyses with different settings.
@@ 4.4 Training Details
$\mathbb{E}_{(\boldsymbol{x},y)\cup D_{n}\sim p(D),\boldsymbol{\lambda}\sim\Delta,g=-s_{\boldsymbol{\lambda}}(\boldsymbol{x})}[-\log q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})].$ (5)
@@
Our final foundation model uses the PFNs architecture with 12 layers of Transformer, an embedding size of 512, a hidden size of 1024 in feed-forward layers, and 4-head attention, resulting in a total of 26.81M parameters. During training, we randomly sample each dataset from a large data space with feature dimensions ranging from 1 to 30, objective dimensions from 1 to 6, and trajectory lengths from 1 to 256. We sample a large variety of functions from GP. Specifically, we randomly sample the length scale kernel parameter $l\sim\Gamma(\alpha=3.0,\beta=6.0)$ for each feature and add a random observation noise $\epsilon\sim\mathcal{N}(0,10^{-4})$ to ensure function diversity. We do not sample the output scale parameter but instead set it to 1.0, as the predicted aggregation target will be normalized during inference. The final model is trained using 500 epochs and 2048 steps in each epoch with a batch size of 128 datasets, resulting more than 130 million datasets in total. The loss function is optimized using Adam optimizer (Kingma and Ba, 2014) with linear warmup and cosine annealing (Loshchilov and Hutter, 2016) using a learning rate of $5\times 10^{-5}$. The entire training process is conducted on 8 NVIDIA L20 GPUs and takes approximately 35 hours. We perform the pre-training only once and use the same model in all experiments in this paper.
@@ 4.5 Generating Candidates
Benefiting from the preference-wise aggregated posterior as the output basis, modeling the distribution of infinite sub-problems becomes possible. Acquisition functions such as random scalarization and multi-objective indicator based on scalarization can be well adapted directly. We have developed two types of acquisition functions including preference-based and preference-free in our work.
Preference-based
We firstly derive the preference-based expected improvement (EI) and upper confidence bound (UCB) acquisition functions based on the single-objective scenario (Müller et al., 2023):
$\alpha_{\text{EI}}(\boldsymbol{x};\boldsymbol{\lambda})=\mathbb{E}_{q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})}[[g_{\boldsymbol{\lambda}}(\boldsymbol{x})-g_{\boldsymbol{\lambda}}^{*}]_{+}],$ (6)
@@
where $g_{\boldsymbol{\lambda}}^{*}$ is the best value of the aggregation $g_{\boldsymbol{\lambda}}(\boldsymbol{x})$ given an arbitrary preference $\boldsymbol{\lambda}$. ${\mu}_{\boldsymbol{\lambda}}(\boldsymbol{x})$ and ${\sigma}_{\boldsymbol{\lambda}}(\boldsymbol{x})$ are the mean and standard deviation of the aggregated posterior $q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})$, respectively. $\beta$ is a constant, which we set to 1.0 in this paper. As inference the acquisition functions is conditioned on the preference, during the optimization process, we randomly sample the preferences in each iteration and optimize the acquisition function to generate the next candidate solutions for evaluation.
Preference-free
To marginalize the preferences, we leverage a key theoretical insight of the popular used hypervolume metric: when employing an infinite set of reference vectors uniformly sampled from the unit hypersphere $S=\{\boldsymbol{w}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{w}\|_{2}=1\}$, the hypervolume metric can be equivalently represented as the mean of a scalarization function (Deng and Zhang, 2019; Zhang and Golovin, 2020). Motivated by this, we firstly develop a preference-free acquisition function that utilizes the UCB utility defined in Equation (7) to approximate the uncertainty-aware hypervolume improvement (UHVI) for any candidate $\boldsymbol{x}$, given the evaluated inputs $\boldsymbol{X}=\{\boldsymbol{x}_{i}\}_{i=1}^{n}$:
@@
where $c_{m}=\frac{\pi^{m/2}}{2^{m}\Gamma(m/2+1)}$ is a constant that depends only on $m$. $c_{\boldsymbol{\lambda}}=\sqrt{\sum_{j=1}^{m}\frac{1}{\lambda_{j}^{2}}}$ is a transformation constant that depends only on $\boldsymbol{\lambda}$. The detailed derivations can be found in Appendix A.3.
Analogous to the derivation of UHVI, we draw upon the R2 indicator (Hansen and Jaszkiewicz, 1994; Brockhoff et al., 2012)—defined as the expected utility across a set of scalarization functions—to develop another preference-free acquisition function. This formulation, termed UR2I (uncertainty-aware R2 improvement), is designed to approximate the R2 improvement while explicitly accounting for predictive uncertainty:
$\alpha_{\text{UR2I}}(\boldsymbol{x})$ $=\mathbb{E}_{\boldsymbol{\lambda}\sim\Delta}[\max\{0,$ (9)
$\min_{\boldsymbol{x}\in\boldsymbol{X}}\{-g_{\boldsymbol{\lambda}}(\boldsymbol{x})\}-[-\alpha_{\text{UCB}}(\boldsymbol{x};\boldsymbol{\lambda})]\}].$
@@ 5 Experiments
We compare our method against a wide range of classic and state-of-the-art MOBO algorithms. For consistency, MOEA/D-EGO (Zhang et al., 2009), TSEMO (Bradford et al., 2018), USEMO-EI (Belakaria et al., 2020), and DGEMO (Konakovic Lukovic et al., 2020) are all implemented using the DGEMO codebase based on Pymoo (Blank and Deb, 2020). In addition, we evaluate qNEHVI (Daulton et al., 2021), qParEGO (Daulton et al., 2020), and JES (Tu et al., 2022; Hvarfner et al., 2022) using the BoTorch framework (Balandat et al., 2020). Finally, we include the recently proposed BOFormer (Hung et al., 2025) as a leaning-based baseline algorithms, given its close relevance to our work.
@@
We first evaluate the performance of all algorithms on 6 synthetic functions featuring a wide spectrum of landscapes, following the experimental setup in BOFormer (Hung et al., 2025). Subsequently, we conduct extensive experiments on 10 real-world engineering design problems from the RE benchmark suite (Tanabe and Ishibuchi, 2020), which encompass diverse decision and objective space dimensions. Finally, we assess the methods on 4 hyperparameter optimization (HPO) tasks in a 3D object reconstruction problem (Mildenhall et al., 2021). Details of test problems are provided in Appendix B.1.
@@
For all evaluated algorithms, each experiment is initialized with $2(d+1)$ randomly generated samples, followed by an optimization budget of 100 evaluations. To assess the quality of the Pareto front approximations produced by different algorithms, we employ the hypervolume (HV) indicator (Zitzler and Thiele, 2002) as the performance metric, given its strict monotonicity with respect to Pareto dominance. All experiments are conducted over 10 independent runs, we report the mean performance alongside the standard deviation with 95% confidence. Further technical details regarding the experimental configurations are provided in Appendix B.2 and B.3.
@@ 5.1 Optimization Performance
Table 1: Means (stds) of hypervolume metrics ($\uparrow$) on synthetic black-box functions over 10 independent runs. The best, second-best, and third-best mean results for each problem are highlighted with dark gray, gray, and light gray backgrounds, respectively.
Algorithm BC AR ABC ABD ACD BCD
MOEA/D-EGO 7.015e-01(7.5e-03) 1.203e+00(2.4e-03) 4.826e-01(8.1e-03) 6.934e-01(1.1e-02) 9.320e-01(1.4e-02) 5.735e-01(1.9e-02)
TSEMO 8.056e-01(9.1e-04) 1.172e+00(1.4e-02) 4.853e-01(4.5e-03) 7.114e-01(9.4e-03) 9.117e-01(2.6e-02) 6.938e-01(1.1e-02)
USEMO-EI 7.750e-01(1.2e-02) 1.118e+00(2.3e-02) 4.648e-01(8.2e-03) 7.298e-01(1.1e-02) 8.991e-01(2.5e-02) 6.766e-01(9.5e-03)
DGEMO 8.087e-01(1.3e-03) 1.158e+00(6.8e-02) 4.477e-01(3.8e-02) 5.918e-01(6.8e-02) 7.266e-01(9.1e-02) 7.054e-01(3.9e-03)
qNEHVI 7.962e-01(1.7e-03) 1.159e+00(1.8e-02) 4.929e-01(6.5e-03) 7.181e-01(1.0e-02) 9.167e-01(1.3e-02) 6.858e-01(1.5e-03)
qParEGO 7.843e-01(4.2e-03) 1.194e+00(4.4e-03) 4.782e-01(7.5e-03) 7.032e-01(9.1e-03) 9.285e-01(1.3e-02) 7.041e-01(4.5e-03)
JES 7.765e-01(3.1e-03) 1.086e+00(3.4e-02) 4.559e-01(1.1e-02) 6.598e-01(1.3e-02) 8.802e-01(2.8e-02) 6.947e-01(3.5e-03)
BOFormer 7.187e-01(1.4e-02) 1.092e+00(3.4e-02) 4.082e-01(1.7e-02) 6.383e-01(2.5e-02) 7.988e-01(3.8e-02) 6.715e-01(1.5e-02)
FoMEMO-EI 8.066e-01(6.2e-04) 1.192e+00(1.0e-02) 4.993e-01(1.2e-02) 7.308e-01(7.9e-03) 9.291e-01(1.5e-02) 7.346e-01(2.3e-03)
FoMEMO-UCB 8.075e-01(5.8e-04) 1.193e+00(8.4e-03) 4.878e-01(1.4e-02) 7.369e-01(8.5e-03) 9.407e-01(1.5e-02) 7.294e-01(6.2e-03)
FoMEMO-UHVI 8.052e-01(4.9e-04) 1.195e+00(6.3e-03) 4.908e-01(7.4e-03) 7.430e-01(8.4e-03) 9.519e-01(1.2e-02) 7.483e-01(7.6e-04)
FoMEMO-UR2I 8.016e-01(1.7e-03) 1.164e+00(2.2e-02) 4.973e-01(6.0e-03) 7.126e-01(9.1e-03) 9.256e-01(1.8e-02) 7.303e-01(1.2e-03)
Table 2: Means (stds) of hypervolume metrics ($\uparrow$) on engineering design problems over 10 independent runs. The best, second-best, and third-best mean results for each problem are highlighted with dark gray, gray, and light gray backgrounds, respectively.
Algorithm Four Bar Truss Pressure Vessel Hatch Cover Disc Brake Speed Reducer Gear Train Rocket Injector
MOEA/D-EGO 7.462e-01(1.4e-02) 1.129e+00(8.8e-03) 1.119e+00(2.3e-02) 1.280e+00(1.4e-02) 1.261e+00(1.3e-02) 5.623e-01(1.3e-01) 7.539e-01(1.1e-02)
TSEMO 8.506e-01(4.7e-04) 1.099e+00(1.6e-02) 1.022e+00(3.5e-02) 1.285e+00(3.0e-02) 1.297e+00(1.3e-03) 8.092e-01(2.3e-02) 8.618e-01(7.5e-04)
USEMO-EI 8.232e-01(4.3e-03) 1.153e+00(1.6e-03) 1.145e+00(6.2e-03) 1.305e+00(2.3e-03) 1.286e+00(3.7e-03) 6.433e-01(4.0e-02) 8.316e-01(3.7e-03)
DGEMO 8.526e-01(9.8e-04) 1.132e+00(6.3e-03) 1.003e+00(7.9e-02) 1.286e+00(2.9e-02) 1.304e+00(1.6e-04) 8.859e-01(3.2e-02) 8.605e-01(1.9e-03)
qNEHVI 8.244e-01(3.8e-03) 1.069e+00(1.6e-02) 1.119e+00(1.3e-02) 1.192e+00(3.1e-02) 1.263e+00(7.9e-03) 8.387e-01(2.6e-02) 8.182e-01(5.0e-03)
qParEGO 7.917e-01(1.7e-02) 1.099e+00(1.5e-02) 1.144e+00(6.7e-03) 1.199e+00(7.4e-02) 1.215e+00(1.4e-02) 7.813e-01(3.1e-02) 6.700e-01(2.8e-02)
JES 7.918e-01(1.0e-02) 8.795e-01(1.3e-01) 1.150e+00(3.5e-03) 1.249e+00(2.5e-02) 1.206e+00(1.0e-02) 6.096e-01(4.5e-02) 6.917e-01(2.4e-02)
BOFormer 7.308e-01(1.4e-02) 3.094e-01(1.7e-01) 1.119e+00(1.5e-02) 1.166e+00(2.3e-02) 1.191e+00(1.0e-02) 3.818e-01(5.7e-02) 6.208e-01(2.2e-02)
FoMEMO-EI 8.667e-01(2.0e-03) 1.151e+00(1.8e-03) 1.162e+00(1.5e-03) 1.309e+00(2.0e-03) 1.298e+00(7.3e-04) 9.107e-01(8.1e-03) 8.506e-01(4.6e-03)
FoMEMO-UCB 8.722e-01(1.2e-03) 1.150e+00(1.8e-03) 1.163e+00(1.9e-03) 1.312e+00(3.1e-04) 1.299e+00(6.2e-04) 9.270e-01(4.5e-03) 8.666e-01(2.4e-03)
FoMEMO-UHVI 8.609e-01(1.3e-03) 1.022e+00(5.4e-02) 1.127e+00(1.1e-02) 1.273e+00(1.6e-02) 1.263e+00(3.7e-03) 8.573e-01(9.6e-03) 8.245e-01(7.3e-03)
FoMEMO-UR2I 8.741e-01(5.0e-04) 1.086e+00(2.3e-02) 1.131e+00(4.8e-03) 1.299e+00(7.2e-03) 1.285e+00(3.6e-03) 8.961e-01(1.1e-02) 8.693e-01(1.5e-03)
Table 3: Means (stds) of hypervolume metrics ($\uparrow$) on hyperparameter optimization problems over 10 independent runs. The best, second-best, and third-best mean results for each problem are highlighted with dark gray, gray, and light gray backgrounds, respectively.
Algorithm Lego Materials Mic Ship
MOEA/D-EGO 1.184e+00(1.1e-03) 1.166e+00(2.2e-03) 1.257e+00(4.9e-03) 1.321e+00(1.5e-03)
TSEMO 1.174e+00(3.3e-03) 1.160e+00(5.1e-03) 1.251e+00(4.8e-03) 1.315e+00(2.1e-03)
USEMO-EI 1.179e+00(1.6e-03) 1.167e+00(2.8e-03) 1.255e+00(5.4e-03) 1.315e+00(3.6e-03)
DGEMO 1.171e+00(5.0e-03) 1.172e+00(3.2e-03) 1.273e+00(2.1e-02) 1.313e+00(3.6e-03)
qNEHVI 1.147e+00(9.8e-03) 1.144e+00(1.1e-02) 1.198e+00(1.9e-02) 1.293e+00(6.6e-03)
qParEGO 1.137e+00(1.2e-02) 1.150e+00(3.3e-03) 1.191e+00(2.0e-02) 1.291e+00(7.6e-03)
JES 1.167e+00(6.6e-03) 1.156e+00(3.2e-03) 1.212e+00(1.3e-02) 1.303e+00(4.5e-03)
BOFormer 1.176e+00(3.1e-03) 1.161e+00(3.6e-03) 1.244e+00(6.3e-03) 1.307e+00(3.3e-03)
FoMEMO-EI 1.177e+00(5.3e-03) 1.173e+00(2.0e-03) 1.261e+00(2.1e-02) 1.315e+00(2.9e-03)
FoMEMO-UCB 1.178e+00(7.3e-03) 1.175e+00(1.9e-03) 1.262e+00(1.8e-02) 1.316e+00(2.3e-03)
FoMEMO-UHVI 1.177e+00(6.2e-03) 1.176e+00(2.4e-03) 1.271e+00(1.6e-02) 1.318e+00(1.9e-03)
FoMEMO-UR2I 1.182e+00(2.6e-03) 1.179e+00(2.2e-03) 1.274e+00(1.2e-02) 1.320e+00(2.1e-03)
Tables 1, 2, and 3 summarize the sequential optimization results across all test problems. The experimental results demonstrate that our framework with four acquisition functions achieves superior performance in the majority of cases. This underscores that by pre-training the foundation model on an extensive and diverse corpus of synthetic data, the model showcases robust adaptability and generalization for problems that it has not encountered during training.
The core intuition driving this advancement lies in the direct and high-fidelity modeling of the unknown aggregated posteriors, combined with the high efficiency of fast in-context optimization. Consequently, it avoids relying on restrictive approximation assumptions of aggregated distributions (like in MOEA/D-EGO) or making excessive trade-offs for computational speed (like in qNEHVI), both of which can introduce modeling inaccuracies and ultimately compromise optimization performance. We further investigate the model’s scalability across varying objective and decision space dimensions, with detailed results and analysis provided in Appendix D.1.